the lenght of a regular lot is 6 yards longer than twice the width.if the perimeter is 288 yards,what are the lots of dimensions

Answer by mananth(16946)

Let the width be x
the length will be 2x+6
2x+2(2x+6)=288
2x+4x+12=288
6x=288-12
6x=276
x=46
length = 2x+6 = 2*46+6 = 98